Three point charges are located it the positive x-axis of a coordinate system. Positive point charge π’’πŸ = 5 𝑛𝐢 and negative point charge π’’πŸ = βˆ’10 𝑛𝐢, are located 3 cm and 5 cm from the origin respectively. Determine the total force of another point charge 𝑸 = βˆ’πŸ 𝝁π‘ͺ at the origin due to the other two chargers.


1
Expert's answer
2022-02-13T17:51:46-0500

"F=\frac{ kq_1q_2}{r^2}"



"F_{1Q}=\frac{9×10^9×5×10^{-9}×1×10^{-6}}{(3×10^{-2})^2}\\=5×10^{-2}N"


"F_{2Q}=\frac{9×10^9×10×10^{-9}×1×10^{-6}}{(5×10^{-2})^2}"


"=3.6\u00d710^{-2}N"


Total Force"=-3.6\u00d710^{-2}+5\u00d710^{-2}"


"=1.4\u00d710^{-2}N"